40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9215     Entry time: Mon Oct 7 14:24:10 2013
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI Config settings re-saved 

I have resaved the PRMI locking settings in the IFO Config screen.  Nothing has changed, except that I have put a 1e-4 into the PRCL matrix elements for REFL11I, REFL33I and REFL55I.  So, PRMI still locks on REFL165 I&Q, but the other 3 REFL diodes' whitening gets triggered when the cavity is locked.  I think this will help the LSC sensing matrix measurements, which I'm going to test out now.

ELOG V3.1.3-